# Stability and convergence of multi-product expansion splitting methods with negative weights for semilinear parabolic equations

### v1

@@ 4 Uniform stability of MPE splitting methods for semilinear parabolic equations
In this section, we derive one of the main results of this work, the uniform stability estimate for semilinear parabolic equations. We introduce the semilinear parabolic equations with periodic boundary conditions. The analysis begins with several lemmas and theorems establishing the stability of the two evolution operators, followed by a complete stability analysis of arbitrarily high-order MPE splitting operators (3.2) which is a highlight of this work. In the following study, we focus on high-order MPE splitting methods with positive time steps but possibly negative weights.
4.1 Semilinear parabolic equations
@@
[…]
where the bounded spatial domain $\Omega=[-\pi,\pi]^{d}$ for $d=1,2,3$ with periodic boundary. The function $f:\mathbb{R}\rightarrow\mathbb{R}$ is nonlinear and $u^{0}(x)$ is the initial data. To establish the stability of $\mathcal{E}_{B}(\tau)$, we impose the following assumption on the function $f$:
@@
$|f^{\prime}|$ is bounded by a constant $\kappa$ on $\mathbb{R}$, i.e., there exists a positive constant $\kappa$ such that
@@
If (1) is not satisfied, i.e., ${\sup}_{\xi\in\mathbb{R}}|f^{\prime}(\xi)|=\infty$, one useful technique is to truncate $f^{\prime}$ when it becomes excessively large so that this assumption can be guaranteed (cf. [41, Eq. (1.8)] for the truncation technique), which will be further discussed in Section 6.
Note that either $f(x)=0$ for some $x\in\mathbb{R}$ or $f(x)\neq 0$ for all $x\in\mathbb{R}$. In the former case, without loss of generality, we assume $f(0)=0$, which is convenient for the proof. In fact, if $f(\xi_{0})=0$ for some constant $\xi_{0}\in\mathbb{R}$, let $w=u-\xi_{0}$. Then, $w$ solves the equation
[…]
where $g(w)=f(w+\xi_{0})$ also satisfies Assumption 1. To recover $u$, we apply a simple translation $u=w+\xi_{0}$, returning to the original problem.
@@
We recall that the evolution operator $\mathcal{E}_{A}(t)$ is instable if $t$ is negative or complex. So, it is natural to assume that the step coefficients $a_{i,j}$ in (3.2) are positive. However, it has been proved theoretically in [42, 46, 22] that negative weights will exist in this case, that is, if $k>2$ and all $a_{i,j}>0$, then $\min_{i}c_{i}<0$. Then, we study the MPE splitting operator defined in (3.2) with all $a_{i,j},b_{i,j}\geq 0$ and finite $m_{i}$. We will later use MPE to refer to MPE splitting methods with all positive step coefficients.
@@ 4.2 Stability of $\mathcal{E}_{A}(\tau)$ and $\mathcal{E}_{B}(\tau)$
    If $u^{0}\in L^{p}(\Omega)$ for $1\leq p<\infty$, then $\|\mathcal{E}_{A}(\tau)u^{0}\|_{L^{p}}\leq\|u^{0}\|_{L^{p}}.$
@@
    If $u^{0}\in L^{\infty}(\Omega)$, then $\|\mathcal{E}_{A}(\tau)u^{0}\|_{L^{\infty}}\leq\|u^{0}\|_{L^{\infty}}.$
@@
(2) Let $u(t)=\mathcal{E}_{A}(t)u^{0}$ for $t\geq 0$. For all convex function $g$,
$\frac{\rm d}{\,{\mathrm{d}}t}\int_{\Omega}g\left(u\right)\,{\mathrm{d}}x=\int_{\Omega}g^{\prime}\left(u\right)\Delta u\,{\mathrm{d}}x=-\int_{\Omega}g^{\prime\prime}\left(u\right)\left|\nabla u\right|^{2}\,{\mathrm{d}}x\leq 0,\quad\forall t\geq 0,$ (4.5)
forms a family of Lyapunov functional. In particular, $g(u)=|u|^{p}$ with $p\in[1,\infty)$ for the $L^{p}$-norm falls into this family.
(3) The linear Laplace operator $\Delta$ is the generator of a contraction semigroup $\{\mathrm{e}^{t\Delta}\}_{t\geq 0}$ and thus naturally possesses the contraction property that $\|\mathcal{E}_{A}(\tau)u^{0}\|_{\infty}\leq\|u^{0}\|_{\infty}$ (see, e.g., [20, Lemma 2.1]). This completes the proof. ∎
The stability analysis of the nonlinear evolution operator $\mathcal{E}_{B}(\tau)$ is lengthier than that of the linear evolution operator $\mathcal{E}_{A}(\tau)$. To facilitate the stability and convergence analysis of the high-order MPE splitting operator, we present the following four lemmas concerning $\mathcal{E}_{B}(\tau)$ and $\mathcal{E}_{B}(\tau)-\mathcal{I}$, where $\mathcal{I}$ denotes the identity operator.
@@
where $\mathcal{I}$ denotes the identity operator and $|\Omega|^{\frac{1}{\infty}}=1$ as aforementioned.
@@
$\left\|[\mathcal{E}_{B}(\tau)-\mathcal{I}]w^{0}\right\|_{p}$ $=\left\|\int^{\tau}_{0}\partial_{t}w\,{\mathrm{d}}t\right\|_{p}=\left\|\int^{\tau}_{0}f(w)\,{\mathrm{d}}t\right\|_{p}$ (4.7)
$\leq\int^{\tau}_{0}\left\|f(w)-f(w^{0})\right\|_{p}\,{\mathrm{d}}t+\int^{\tau}_{0}\left\|f(w^{0})\right\|_{p}\,{\mathrm{d}}t$
$\leq\kappa\int^{\tau}_{0}\left\|w-w^{0}\right\|_{p}\,{\mathrm{d}}t+\tau\left\|f(w^{0})\right\|_{p}$
$=\kappa\int^{\tau}_{0}\left\|[\mathcal{E}_{B}(t)-\mathcal{I}]w^{0}\right\|_{p}\,{\mathrm{d}}t+\tau\left\|f(w^{0})\right\|_{p}.$
Applying Grönwall’s lemma, we obtain
@@
$\leq\int^{\tau}_{0}\left|f(w)-f(v)\right|\,{\mathrm{d}}t+\left|w^{0}(x)-v^{0}(x)\right|$
$\leq\kappa\int^{\tau}_{0}|w-v|\,{\mathrm{d}}t+\left|w^{0}(x)-v^{0}(x)\right|.$
@@
[…]
@@
4.3 Stability of genaral MPE splitting methods
@@
and
$\left\|\mathcal{S}(\tau)w^{0}-\mathcal{E}_{A}(\widetilde{a}\tau)w^{0}\right\|_{p}\leq\left(\mathrm{e}^{\kappa\widetilde{b}\tau}-1\right)\left(\left\|w^{0}\right\|_{p}+\frac{|\Omega|^{\frac{1}{p}}|f(0)|}{\kappa}\right),$ (4.19)
@@
${L}_{1}$ $\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\left\|\mathcal{E}_{A}(a_{m}\tau)\left[\widetilde{\mathcal{S}}(\tau)w^{0}-\widetilde{\mathcal{S}}(\tau)v^{0}\right]\right\|_{p}$ (4.23)
$\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\left\|\widetilde{\mathcal{S}}(\tau)w^{0}-\widetilde{\mathcal{S}}(\tau)v^{0}\right\|_{p}$
$\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\mathrm{e}^{\kappa\widetilde{b}_{m-1}\tau}\left\|w^{0}-v^{0}\right\|_{p}$
$=\kappa b_{m}\mathrm{e}^{\kappa\widetilde{b}_{m}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p},$
@@
$\leq$ $\kappa(b_{m}+b_{m-1}+\cdots+b_{1})\mathrm{e}^{\kappa\widetilde{b}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p}$
$=$ $\kappa\widetilde{b}\mathrm{e}^{\kappa\widetilde{b}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p}.$
@@
where $C=(\widetilde{c}+1)\kappa b$, $\widetilde{c}=\sum^{M}_{i=1}|c_{i}|$, and $b=\max_{1\leq i\leq M}\left\{\sum_{j=1}^{m_{i}}b_{i,j}\right\}$.
Moreover, for any $T>0$ and $0<n\tau\leq T$, the uniform bounds of $\mathcal{S}^{[k]}(\tau)$ can be derived as
@@
It should be noted that the order condition (3.14) inherently enforces the condition $\sum_{i=1}^{M}c_{i}=1$ for any $k$th-order ($k\geq 1$) MPE splitting operator, while the condition $\sum_{j=1}^{m_{i}}a_{i,j}=1$ for $1\leq i\leq M$ is imposed to obtain the stability result in Theorem 4.1.
@@
$\left\|\mathcal{S}^{[k]}(\tau)w^{0}-\mathcal{S}^{[k]}(\tau)v^{0}\right\|_{p}$ $\leq(\widetilde{c}\kappa b\mathrm{e}^{\kappa b\tau}\tau+1)\left\|w^{0}-v^{0}\right\|_{p}$ (4.31)
$\leq\mathrm{e}^{(\widetilde{c}+1)\kappa b\tau}\left\|w^{0}-v^{0}\right\|_{p}.$
@@
$\leq\sum^{M}_{i=1}|c_{i}|\left(\mathrm{e}^{\kappa\widetilde{b}_{i,m_{i}}\tau}-1\right)\left(\left\|w^{0}\right\|_{p}+\frac{\eta}{\kappa}\right)+\left\|w^{0}\right\|_{p}$
$\leq\mathrm{e}^{\widetilde{c}\kappa b\tau}\left\|w^{0}\right\|_{p}+\left(\mathrm{e}^{\widetilde{c}\kappa b\tau}-1\right)\frac{\eta}{\kappa}.$
Therefore, we can establish uniform bounds on the numerical solution $u^{n}$ at $t_{n}$-level as follows:
@@ 5 Convergence of MPE splitting methods for semilinear parabolic equations
In this section, a rigorous local error analysis and convergence of MPE splitting methods for semilinear parabolic equations are given. We first derive the uniform bounds of the exact solution to the semilinear parabolic equations. To derive the global temporal error estimate, we employ the formal calculus of Lie derivatives and iterated Lie-commutators, which help to establish the consistency of MPE splitting methods. Finally, we establish the convergence results of MPE splitting methods based on the previous stability results.

### v2

@@ 4 Stability of MPE splitting methods for semilinear parabolic equations
In this section, we derive one of the main results of this work, the stability estimate for semilinear parabolic equations. We introduce the semilinear parabolic equations with periodic boundary conditions. The analysis begins with several lemmas and theorems establishing the stability of the two evolution operators, followed by a complete stability analysis of arbitrarily high-order MPE splitting operators (3.2) which is a highlight of this work. In the following study, we focus on high-order MPE splitting methods with positive time steps but possibly negative weights.
4.1 Semilinear parabolic equation
@@
$\left\{\begin{aligned} &u_{t}=\Delta u+f(u),&&x\in\Omega,~t\in(0,T],\\ &u(x,0)=u^{0}({x}),&&x\in\Omega,\end{aligned}\right$. (4.1)
where the bounded spatial domain $\Omega=[-\pi,\pi]^{d}$ for $d=1,2,3$. The function $f:\mathbb{R}\rightarrow\mathbb{R}$ is nonlinear and $u^{0}(x)$ is the initial data. To establish the stability of $\mathcal{E}_{B}(\tau)$, we impose the following assumption on the function $f$:
@@
There exists a positive constant $\kappa$ such that
@@
Remark 4.1.
From the existence and uniqueness theorem for nonlinear ODEs, if $f$ is only locally Lipschitz, the solutions may not be defined for certain values of $t$, even if $f$ is smooth. For instance, the differential equation $u_{t}=u^{2}$ with initial condition $u(0)=1$ has the solution u(t)=1/(1t)u(t)=1/(1-t), which is not defined at $t=1$. This implies that operator splitting methods can never be stable unconditionally. Therefore, Assumption 1 is reasonable.
Remark 4.2.
If (1) is not satisfied, one technique is to truncate $f^{\prime}$ when it becomes excessively large so that this assumption can be guaranteed (cf. [41, Eq. (1.8)] for the truncation technique), which will be further discussed in Section 6.
@@
For splitting methods with real step coefficients, we recall that the evolution operator $\mathcal{E}_{A}(t)$ is unstable if $t$ is negative. So, it is natural to assume that the step coefficients $a_{i,j}$ in (3.2) are positive. However, it has been proved theoretically in [42, 46, 25] that negative weights will exist in this case, that is, if $k>2$ and all $a_{i,j}>0$, then $\min_{i}c_{i}<0$. Then, we study the MPE splitting operator defined in (3.2) with all $a_{i,j},b_{i,j}\geq 0$ and finite $m_{i}$. We will later use MPE to refer to MPE splitting methods with all positive step coefficients.
@@ 4.2 Stability of $\mathcal{E}_{A}(\tau)$ and $\mathcal{E}_{B}(\tau)$
    If $u^{0}\in L^{p}(\Omega)$ for $1\leq p<\infty$, then $\|\mathcal{E}_{A}(\tau)u^{0}\|_{p}\leq\|u^{0}\|_{p}.$
@@
    If $u^{0}\in L^{\infty}(\Omega)$, then $\|\mathcal{E}_{A}(\tau)u^{0}\|_{\infty}\leq\|u^{0}\|_{\infty}.$
@@
(2) Let $u(t)=\mathcal{E}_{A}(t)u^{0}$ for $t\geq 0$. Consider a sequence of smooth convex functions $g_{\varepsilon}(u)=(u^{2}+\varepsilon^{2})^{p/2}$ with small parameter $\varepsilon\in(0,1]$. Straightforward calculations give
$\frac{\rm d}{\,{\mathrm{d}}t}\int_{\Omega}g_{\varepsilon}\left(u\right)\,{\mathrm{d}}x=\int_{\Omega}g_{\varepsilon}^{\prime}\left(u\right)\Delta u\,{\mathrm{d}}x=-\int_{\Omega}g_{\varepsilon}^{\prime\prime}\left(u\right)\left|\nabla u\right|^{2}\,{\mathrm{d}}x\leq 0,\qquad\forall t\geq 0.$ (4.4)
Then, for any $\varepsilon\in(0,1]$, we have
$\int_{\Omega}|u(t)|^{p}\,{\mathrm{d}}x\leq\int_{\Omega}g_{\varepsilon}(u(t))\,{\mathrm{d}}x\leq\int_{\Omega}g_{\varepsilon}(u^{0})\,{\mathrm{d}}x.$ (4.5)
Since $\lim_{\varepsilon\rightarrow 0}g_{\varepsilon}(u^{0})=|u^{0}|^{p}$ and $g_{\varepsilon}(u^{0})\leq((u^{0})^{2}+1)^{p/2}$, we can directly obtain
$\lim_{\varepsilon\rightarrow 0}\int_{\Omega}g_{\varepsilon}(u^{0})\,{\mathrm{d}}x=\int_{\Omega}|u^{0}|^{p}\,{\mathrm{d}}x$ (4.6)
by applying the dominated convergence theorem. As a consequence, taking $\varepsilon\rightarrow 0$ in (4.5), we then have the desired result.
(3) The linear Laplace operator $\Delta$ is the generator of a contraction semigroup $\{\mathrm{e}^{t\Delta}\}_{t\geq 0}$ and thus naturally possesses the contraction property that $\|\mathcal{E}_{A}(\tau)u^{0}\|_{\infty}\leq\|u^{0}\|_{\infty}$ (see, e.g., [21, Lemma 2.1]). This completes the proof. ∎
In fact, the $H^{k}$-norm contractivity in Lemma 4.1 will be used in the subsequent consistency analysis. The stability analysis of the nonlinear evolution operator $\mathcal{E}_{B}(\tau)$ is lengthier than that of the linear evolution operator $\mathcal{E}_{A}(\tau)$. To facilitate the stability and convergence analysis of the high-order MPE splitting operator, we present the following four lemmas concerning $\mathcal{E}_{B}(\tau)$ and $\mathcal{E}_{B}(\tau)$-$\mathcal{I}$, where $\mathcal{I}$ denotes the identity operator.
@@
where $\mathcal{I}$ denotes the identity operator, $|\Omega|$ is the volume of the bounded domain $\Omega$, and $|\Omega|^{\frac{1}{\infty}}=1$ as aforementioned.
@@
$\left\|[\mathcal{E}_{B}(\tau)-\mathcal{I}]w^{0}\right\|_{p}$ $=\left\|\int^{\tau}_{0}\partial_{t}w\,{\mathrm{d}}t\right\|_{p}=\left\|\int^{\tau}_{0}f(w)\,{\mathrm{d}}t\right\|_{p}\leq\int^{\tau}_{0}\left\|f(w)-f(w^{0})\right\|_{p}\,{\mathrm{d}}t+\int^{\tau}_{0}\left\|f(w^{0})\right\|_{p}\,{\mathrm{d}}t$ (4.8)
$\leq\kappa\int^{\tau}_{0}\left\|w-w^{0}\right\|_{p}\,{\mathrm{d}}t+\tau\left\|f(w^{0})\right\|_{p}=\kappa\int^{\tau}_{0}\left\|[\mathcal{E}_{B}(t)-\mathcal{I}]w^{0}\right\|_{p}\,{\mathrm{d}}t+\tau\left\|f(w^{0})\right\|_{p},$
where $\kappa$ is introduced in (4.2). Applying Grönwall’s lemma, we obtain
@@
$\leq\int^{\tau}_{0}\left|f(w)-f(v)\right|\,{\mathrm{d}}t+\left|w^{0}(x)-v^{0}(x)\right|\leq\kappa\int^{\tau}_{0}|w-v|\,{\mathrm{d}}t+\left|w^{0}(x)-v^{0}(x)\right|.$
@@
$\left|[\mathcal{E}_{B}(\tau)-\mathcal{I}]w^{0}(x)-[\mathcal{E}_{B}(\tau)-\mathcal{I}]v^{0}(x)\right|$ $=\left|\int_{0}^{\tau}\left(f(w)-f(v)\right)dt\right|$ (4.18)
@@
4.3 Stability of general MPE splitting methods
@@
$\left\|\mathcal{S}(\tau)w^{0}-\mathcal{E}_{A}(\widetilde{a}\tau)w^{0}\right\|_{p}\leq\left(\mathrm{e}^{\kappa\widetilde{b}\tau}-1\right)\left(\left\|w^{0}\right\|_{p}+\frac{|\Omega|^{\frac{1}{p}}|f(0)|}{\kappa}\right),$
@@
${L}_{1}$ $\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\left\|\mathcal{E}_{A}(a_{m}\tau)\left[\widetilde{\mathcal{S}}(\tau)w^{0}-\widetilde{\mathcal{S}}(\tau)v^{0}\right]\right\|_{p}\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\left\|\widetilde{\mathcal{S}}(\tau)w^{0}-\widetilde{\mathcal{S}}(\tau)v^{0}\right\|_{p}$ (4.23)
$\leq\kappa b_{m}\mathrm{e}^{\kappa b_{m}\tau}\tau\mathrm{e}^{\kappa\widetilde{b}_{m-1}\tau}\left\|w^{0}-v^{0}\right\|_{p}=\kappa b_{m}\mathrm{e}^{\kappa\widetilde{b}_{m}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p},$
@@
$\leq$ $\kappa(b_{m}+b_{m-1}+\cdots+b_{1})\mathrm{e}^{\kappa\widetilde{b}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p}=\kappa\widetilde{b}\mathrm{e}^{\kappa\widetilde{b}\tau}\tau\left\|w^{0}-v^{0}\right\|_{p}.$
@@
where $C=(\widetilde{c}+1)\kappa b$, $\widetilde{c}=\sum^{M}_{i=1}|c_{i}|$, and $b=\max_{1\leq i\leq M}\left\{\sum_{j=1}^{m_{i}}b_{i,j}\right\}$. Moreover, for any $T>0$ and $0<n\tau\leq T$, the Sobolev bounds of $\mathcal{S}^{[k]}(\tau)$ can be derived as
@@
It should be noted that the order condition of MPE splitting inherently enforces the condition $\sum_{i=1}^{M}c_{i}=1$ for any $k$th-order ($k\geq 1$) MPE splitting operator, while the condition $\sum_{j=1}^{m_{i}}a_{i,j}=1$ for $1\leq i\leq M$ is imposed to obtain the stability result in Theorem 4.1.
@@
$\left\|\mathcal{S}^{[k]}(\tau)w^{0}-\mathcal{S}^{[k]}(\tau)v^{0}\right\|_{p}$ $\leq(\widetilde{c}\kappa b\mathrm{e}^{\kappa b\tau}\tau+1)\left\|w^{0}-v^{0}\right\|_{p}\leq\mathrm{e}^{(\widetilde{c}+1)\kappa b\tau}\left\|w^{0}-v^{0}\right\|_{p}.$ (4.31)
@@
$\leq\sum^{M}_{i=1}|c_{i}|\left(\mathrm{e}^{\kappa\widetilde{b}_{i,m_{i}}\tau}-1\right)\left(\left\|w^{0}\right\|_{p}+\frac{\eta}{\kappa}\right)+\left\|w^{0}\right\|_{p}\leq\mathrm{e}^{\widetilde{c}\kappa b\tau}\left\|w^{0}\right\|_{p}+\left(\mathrm{e}^{\widetilde{c}\kappa b\tau}-1\right)\frac{\eta}{\kappa}.$
Therefore, we can establish Sobolev bounds on the numerical solution $u^{n}$ at $t_{n}$-level as follows:
@@
Remark 4.4.
The above stability analysis can also be extended to homogeneous Neumann boundary conditions, as the stability of the solution operators to both subproblems. However, for general Dirichlet, Robin, and Neumann boundary conditions, order reduction might happen even for Strang splitting (see for example [22, 23, 4]). It is still a challenging problem to investigate the stability of MPE splittings for these boundary conditions.
@@ 5 Convergence of MPE splitting methods for semilinear parabolic equations
In this section, a rigorous local error analysis and convergence of MPE splitting methods for semilinear parabolic equations are given. We first derive the Sobolev bounds of the exact solution to the semilinear parabolic equations. To derive the global temporal error estimate, we employ the formal calculus of Lie derivative and iterated Lie-commutators, which help to establish the consistency of MPE splitting methods. Finally, we establish the convergence results of MPE splitting methods based on the previous stability results.